# Face Reconstruction Transfer Attack as Out-of-Distribution Generalization

### v1

@@ 4.4 Comparison with Previous Works
We compare our method with state-of-the-art feature-based face reconstruction methods including NBNet [18], LatentMap [6], Genetic [7], GaussBlob [23], Eigenface [24], FaceTI [27], and QEZOGE [19]. For FaceTI [27], we used StyleGAN2 and our face encoders for reproduction. As shown in Tab. 1 and Tab. 2, overall previous methods are effective on seen encoders, but the performance drastically drops on unseen encoders. EigenFace [24], FaceTI [27], and QEZOGE [19] show better performance compared with other works, however, tend to show lower performance compared with our method and the results highly fluctuates depending on the type of seen encoder. In contrast, our method outperforms for both seen and unseen cases. Our SAR and identification rate results are close to real face images on seen encoders while outperforming previous works with a large margin on unseen encoders for every dataset while depending less on the type of seen encoder. Additionally, results tested on unseen encoders shown in Tab. 1 and Tab. 2 present that our method achieves OOD generalization on unseen encoders successfully.
Table 3: Ablation study controlling the components of our suggested method. Performance is averaged over the test results on unseen encoders with SAR@FAR$(=1e-4,1e-3,1e-2)$ and SAR$(score\geq threshold)$ for each of the datasets respectively. Best results are highlighted in bold.
@@ 4.5.1 Ablation of Components
We analyze the effect of each component of our method on the overall performance. In Tab. 3, we present results for $n=1,20,50,100$ with and without latent averaging and unsupervised validation. We use $k_{top}=10$ and $t=70$ for default when applied. In addition to SAR, we report SAR@FAR(=1e-4, 1e-3, 1-e2) scores which are similar to the concept of TAR@FAR widely used in T/F evaluations which facilitates a more thorough analysis of our work. Results in Tab. 3 reveal that using more initial samples and applying ALSUV yields the best performance. Compared to optimizing a single latent without ALSUV, our method increases SAR by 46.14% for LFW, 12.23% for CFP-FP, and 22.62% for AgeDB-30 on average.
@@ 4.5.2 Analysis of Hyperparameters
In this section, we investigate the effects of each hyperparameter related to our method. Our method has 3 parameters: number of latents $n$, size of latent average $t$, and number of top $k$ samples for unsupervised validation $k_{top}$. We evaluate results by controlling these hyperparameters tested on unseen encoders on LFW dataset.
First of all, we consider $1\leq n\leq 100$ with (blue) and without (red) ALSUV. As shown in Fig. 3a, the $n$ plays a crucial role as SAR increases from 48.66% to 92.06% (red line). SAR increases remarkably, especially within the range of $1\leq n\leq 10$, and starts plateauing from $20\leq n$. In addition, we can observe the evident influence of latent averaging and unsupervised validation from the gap between the two lines. Finally, improvement starts plateauing from $n\geq 20$ as shown in Fig. 3a and Tab. 3. Therefore, we suggest $n=20$ as a cost-efficient trade-off point between high performance and computation.
Secondly, we compare the effect of the size of latent average $t$ for $t\in$1, 10, 20, 30, 40, 50, 60, 70, 80, 90, 100} where we fix $n=100$ and test with and without unsupervised validation. As shown in Fig. 3b, the overall performance is highest in the interval $70\leq t\leq 90$. Compared with $t=1$ which is without latent averaging, latent averaging improves the performance from 92.06% to 93.46% without unsupervised validation, and 94.24% to 95.13% with unsupervised validation. Interestingly, we found that any size of latent averaging benefits when unsupervised validation is not applied(red line).
@@ 4.5.3 Number of Latents and Optimization Steps
We compare optimizing 100 latents optimized 100 steps each (denoted as 100 $\times$ 100, without latent average and unsupervised validation) and a single latent optimized 10,000 steps with cyclical learning rate[13]. We train on LFW dataset where results are shown in Tab. 6. Our method outperforms with 95.13% in average SAR while serial optimization shows 60.28% resulting in a 34.85% performance gap. Despite the total steps of optimizations being identical, results highlight the importance of using multiple latents which prevents falling into poor local minima than interacting with complex learning rate scheduler as the difference of average cosine similarity shown in the last row of Tab. 6 is significant.
@@ 4.5.5 Unsupervised Validation with Pseudo Target
Unsupervised validation utilizes the feature space of a surrogate validation encoder to search for a better generalizing sample instead of only using the seen encoder. Hence, we compare the distance between the validation encoder’s top 1 and seen encoder’s top 1 against the ground truth feature in validation space (Fig. 5a) and examine whether this is relevant to generalizing to unseen encoder space (Fig. 5b). We also measure the distance between pseudo targets against ground truth in validation space to verify its efficacy as a target. We use LFW dataset and examine all 6 encoders as seen and unseen between each other as Tab. 1 with Swin-T as validation encoder and acquire the statistics of cosine distance.
As shown in Fig. 5a the pseudo target is closest to the ground truth feature from real images with average cosine distance $0.473$ followed by our unsupervised validation’s top 1 with $0.386$ and seen encoder’s top1 with $0.309$ in the validation space. According to this result, the pseudo target which is the closest to ground truth might be the best option, but unfortunately, the corresponding latent is inaccessible which is why we use the sample closest to the pseudo target. This aspect is connected to the unseen encoder space in Fig. 5b where our method shows average cosine distance $0.29$ higher than seen encoder’s top 1 $0.241$.
Furthermore, we examine results by changing the type of validation encoder. We use each encoder used in our work as a validation encoder and examine the performance improvement only for unseen cases where seen, unseen, and validation encoders do not overlap. Results shown in Fig. 5c presents that using any type of validation encoder improves performance and the improvement shows positive correlation with the general performance of validation encoder. More results are shown in Supp. 0.C Tab. 7.
Figure 5: Statistics of cosine distance between seen encoder’s top 1, unsupervised validation’s top 1, and pseudo target against real image’s feature in validation feature space (a) and unseen encoder space (b) (pseudo target excluded since it only presents in validation space). (c) shows the correlation graph between the validation encoder’s performance and performance improvement.
@@ 4.5.6 Image Quality
We conduct qualitative and quantitative analyses of generated images. For quality evaluation, we compare a few samples from LFW and AgeDB in Fig. 6 and CFP-FP in Supp. 0.C Fig. 7. For quantitative evaluation, we adopt face-specific quality metrics SER-FIQ [30] and CR-FIQA [2]. NBNet, GaussBlob, and EigenFace show poor image quality visually with artifacts and quantitatively low image quality metrics. Meanwhile, despite the decent quality, LatentMap, Genetic, and FaceTI present wrong identities compared with the target images. On the other hand, QEZOGE and our method show decent image quality visually, quantitatively, and content-wise as shown in Tab. 6.
@@ 5 Conclusion
In this paper, we have presented a framework for face reconstruction transfer attacks. We devised our method inspired by out-of-distribution generalization to generalize our generated sample to unseen face encoders and propose ALSUV. ALSUV is instantiated by combining multiple latent optimization, latent averaging, and unsupervised validation with the pseudo target. We demonstrate that our approach surpasses previous methods in FRTA by showing high SAR and identification rate across various unseen face encoders. Our thorough analysis shows the effectiveness of our method inspired by OOD generalization. Furthermore, we hope our work alerts the security risk posed by FRTA, and emphasizes the awareness to mitigate potential threats.
@@ Appendix 0.C Supplementary Results
Tab. 7 ablates pseudo target of unsupervised validation by using different types of target for searching the top 1 reconstructed sample. We compare SAR of 3 different cases: using the seen feature vector as target in the seen encoder space, using validation encoder and the pseudo target in the validation encoder space, and using the unseen encoder and the feature vector from real image in the unseen encoder space where the last works as a reference to upper bound performance.
Table 7: SAR comparison by using different targets when choosing top 1 sample. We compare using seen feature vector in seen encoder space(equivalent to not using unsupervised validation) denoted as Seen Feature, our pseudo target in validation encoder space denoted as Pseudo Target, and feature vector of real image in unseen encoder space which works as reference to upper bound which is denoted as Upper Bound. Unsupervised validation results are highlighted with gray and best performances (except upper bound) are highlighted with bold.

### v2

@@ 4.4 Comparison with Previous Works
We compare our method with state-of-the-art feature-based face reconstruction methods including NBNet [21], LatentMap [7], Genetic [8], GaussBlob [26], Eigenface [27], FaceTI [32], and QEZOGE [22]. For FaceTI [32], we used StyleGAN2 and our face encoders for reproduction. As shown in Tab. 1 and Tab. 2, overall previous methods are effective on seen encoders, but the performance drastically drops on unseen encoders. EigenFace [27], FaceTI [32], and QEZOGE [22] show better performance compared with other works, however, tend to show lower performance compared with our method and the results highly fluctuates depending on the type of seen encoder. In contrast, our method outperforms for both seen and unseen cases. Our SAR and identification rate results are close to real face images on seen encoders while outperforming previous works with a large margin on unseen encoders for every dataset while depending less on the type of seen encoder. Additionally, results tested on unseen encoders shown in Tab. 1 and Tab. 2 present that our method achieves OOD generalization on unseen encoders successfully. We additionally attack SOTA face recognition systems AdaFace [17], CurricularFace [14], and ElasticFace [2] in Tab. 8.
Table 3: Ablation study controlling the components of our method. Performance is averaged over the test results on unseen encoders with SAR@FAR$(=1e-4,1e-3,1e-2)$ and SAR$(score{\geq}threshold)$ for each of the datasets respectively. Best results are highlighted in bold.
@@ 4.5.1 Ablation of Components
We analyze the effect of each component of our method. In Tab. 3, we present results for $n=1,20,50,100$ with and without latent averaging and unsupervised validation. We use $k_{top}=10$ and $t_{avg}=70$ for default when applied. We additionally report SAR@FAR(=1e-4, 1e-3, 1-e2), which is the success attack rate at thresholds of FAR. SAR@FAR is similar to TAR@FAR widely used in T/F evaluations, facilitating a more thorough analysis. Results in Tab. 3 reveal that using more initial samples and applying ALSUV yield the best performance. Comparing with optimizing a single latent without ALSUV, our method increases SAR by 46.14% for LFW, 12.23% for CFP-FP, and 22.62% for AgeDB-30 on average.
@@ 4.5.2 Analysis of Hyperparameters
We investigate the effects of each hyperparameter: number of latents $n$, size of latent average $t_{avg}$, and number of top $k$ samples for unsupervised validation $k_{top}$. We evaluate results by controlling these hyperparameters tested on unseen encoders on LFW dataset.
First of all, we consider $1\leq n\leq 100$ with (blue) and without (red) ALSUV. As shown in Fig. 3a, $n$ plays a crucial role as SAR increases from 48.66% to 92.06% (red line). SAR increases remarkably, especially within the range of $1\leq n\leq 10$, and starts plateauing from $20\leq n$ as shown in Fig. 3a and Tab. 3. Therefore, we suggest $n=20$ as a cost-efficient trade-off point between high performance and computation. Additionally, the gap between red and blue lines shows the influence of latent averaging and unsupervised validation.
Next, we analyze the effect of the size of latent average $t_{avg}$ for $t_{avg}\in$1, 10, 20, 30, 40, 50, 60, 70, 80, 90, 100} with $n=100$ and test with and without unsupervised validation. As shown in Fig. 3b, the overall performance is highest for $70\leq t_{avg}\leq 90$. Compared with $t_{avg}=1$ (without latent averaging), latent averaging improves the performance from 92.06% to 93.46% without unsupervised validation, and 94.24% to 95.13% with unsupervised validation. Interestingly, any size of latent averaging benefits without unsupervised validation(red line).
@@ 4.5.3 Number of Latents and Optimization Steps
We compare optimizing 100 latents optimized 100 steps each (denoted as 100 $\times$ 100, without latent average and unsupervised validation) and a single latent optimized 10,000 steps with cyclical learning rate[15]. We train on LFW dataset where results are shown in Tab. 6. Our method outperforms with 95.13% in average SAR while serial optimization shows 60.28% resulting in a 34.85% performance gap. Despite the total steps of optimization being identical, results highlight the importance of using multiple latents which prevents falling into poor local minima than interacting with complex learning rate scheduler as the difference of average cosine similarity shown in the last row of Tab. 6 is significant.
@@ 4.5.5 Unsupervised Validation with Pseudo Target
Unsupervised validation utilizes the feature space of a surrogate validation encoder to search for a better generalizing sample instead of only searching the seen encoder feature space. Hence, we compare the distance between the validation encoder’s top 1 and seen encoder’s top 1 against the ground truth feature in validation space (Fig. 5a) and examine whether this is relevant to generalizing to unseen encoder space (Fig. 5b). We also measure the distance between pseudo targets against ground truth in validation space to verify its efficacy as a target. We use LFW dataset and examine all 6 encoders as seen and unseen between each other as Tab. 1 with Swin-T as validation encoder.
As shown in Fig. 5a the pseudo target is closest to the ground truth feature from real images with average cosine distance $0.473$ followed by our unsupervised validation’s top 1 with $0.386$ and seen encoder’s top 1 with $0.309$ in the validation space. The closest pseudo target might be the best option to use, but unfortunately, its corresponding latent is inaccessible which makes the sample closest to the pseudo target our best choice. This aspect is connected to the unseen encoder space in Fig. 5b where our method shows average cosine distance $0.29$ higher than seen encoder’s top 1 $0.241$.
Furthermore, we examine performances by varying the type of validation encoder. Each encoders used in our work are used as validation encoders and performance improvement only for unseen cases where seen, unseen, and validation encoders do not overlap are measured. Results in Fig. 5c presents that using any validation encoder shows improvement with positive correlation with the general performance of the encoder. More results are shown in Supp. 0.C.1 Tab. 7.
Figure 5: Cosine distance statistics between seen encoder’s top 1, unsupervised validation’s top 1, and pseudo target against real image’s feature in validation feature space (a) and unseen encoder space (b) (pseudo target excluded since it only presents in validation space). (c) shows the correlation between the validation encoder’s performance and performance improvement. 88
@@ 4.5.6 Image Quality
We conduct qualitative and quantitative analyses of generated images. For quality evaluation, we compare a few samples from LFW and AgeDB in Fig. 6 and CFP-FP in Supp. 0.C.1 Fig. 7. For quantitative evaluation, we adopt face-specific quality metrics SER-FIQ [35] and CR-FIQA [3]. NBNet, GaussBlob, and EigenFace show poor image quality visually with artifacts and quantitatively low image quality metrics. Meanwhile, despite the decent quality, LatentMap, Genetic, and FaceTI present wrong identities compared with the target images. On the other hand, QEZOGE and our method show decent image quality visually, quantitatively, and content-wise as shown in Tab. 6.
@@ 5 Conclusion
In this paper, we have presented a framework for face reconstruction transfer attacks. We devised our method ALSUV inspired by out-of-distribution generalization to generalize our generated sample to unseen face encoders. ALSUV is instantiated by combining multiple latent optimization, latent averaging, and unsupervised validation with the pseudo target. We demonstrate that our approach surpasses previous methods in FRTA by showing high SAR and identification rate across various unseen face encoders. Our thorough analysis shows the effectiveness of our method inspired by OOD generalization. Furthermore, we hope our work alerts the security risk posed by FRTA, and emphasizes the awareness to mitigate potential threats.
Acknowledgements
This work is supported in part by NSF grant CNS-2038493, ONR grant N00014-21-1-2431 from NCI, and U.S. Department of Homeland Security grant 22STESE00001-03-02. The views and conclusions contained in this document are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of the U.S. Department of Homeland Security, the NSF and the ONR.
@@ 0.C.1 Sample Images
0.C.2 Unsupervised Validation Analysis
Tab. 7 ablates pseudo target of unsupervised validation by using different types of target for searching the top 1 reconstructed sample. We compare 3 different cases: using the seen feature as target in the seen encoder space, using validation encoder and the pseudo target in the validation encoder space, and using the unseen encoder and the feature from real image in the unseen encoder space where the last works as a reference to upper bound performance.
Table 7: SAR comparison by using different targets when choosing top 1 sample. Unsupervised validation results are highlighted with gray and best performances (except upper bound) are highlighted with bold.
@@ 0.C.2 Unsupervised Validation Analysis
0.C.3 Attacking SOTA Face Recognition Systems
Table 8: SAR of previous works and our proposed method with SOTA target encoders: AdaFace [17], CurricularFace [14], and ElasticFace [2]. We highlight our method with grey and best results are highlighted with bold.
Dataset LFW
Test Encoder FaceNet MobFace ResNet50 ResNet100 Swin-S VGGNet Unseen AVG
Target Encoder Real Face 99.87 99.83 99.83 99.93 99.93 99.8 -
AdaFace NBNet 49.27 12.99 61.56 20.18 36.31 41.92 37.04
LatentMap 15.36 2.06 13.4 3.88 1.55 9.15 7.57
Genetic 14.17 3.58 7.19 7.49 12.22 8.23 8.81
GaussBlob 0.03 30.14 0.07 85.59 10.29 0.03 21.03
EigenFace 32.97 80.16 68.51 98.31 66.86 77.49 70.72
FaceTI 0.67 7.93 3.24 26.49 52.55 0.54 15.24
QEZOGE 73.41 63.75 72.7 83.8 83.53 74.65 75.31
Ours 88.3 96.8 97.37 99.33 99.09 95.75 96.11
CurricularFace NBNet 58.72 10.53 63.99 58.18 39.79 46.81 46.34
LatentMap 28.99 5.26 19.14 9.52 4.22 14.38 13.59
Genetic 31.69 8.2 17.25 21.53 23.96 19.27 20.32
GaussBlob 0.64 27.98 0.51 86.6 6.65 0.37 20.46
EigenFace 51.54 79.95 79.24 97.91 56.06 82.52 74.54
FaceTI 2.36 16.06 9.58 47.65 58.42 1.92 22.67
QEZOGE 86.2 67.09 80.97 90.08 84.75 82.52 81.94
Ours 93.97 95.99 96.73 99.29 98.38 95.99 96.73
ElasticFace NBNet 50.86 6.78 65.47 57.34 39.66 45.87 44.33
LatentMap 27 5.13 16.94 8.98 4.32 14.14 12.75
Genetic 30.81 7.9 14.88 23.08 19.44 15.86 18.66
GaussBlob 0.03 12.42 0.34 84.81 5.16 0.17 17.16
EigenFace 52.28 71.25 75.13 96.69 54.88 81.3 71.92
FaceTI 3.1 15.52 9.75 53.39 59.4 2.26 23.90
QEZOGE 82.79 66.66 81.4 92.2 84.58 81.07 81.45
Ours 94.24 95.75 97.03 99.29 98.21 96.49 96.84
Additionally, we validate the efficacy of ALSUV by attacking SOTA face recognition systems: AdaFace [17], CurricularFace [14], and Elastic [2] with the samples generated from the 6 encoders from Tab. 1 and LFW dataset. ALSUV is compared with previous methods and all resuls are unseen scenario as all 6 encoders do not overlap with the additional face recognition systems. Results are shown in Tab. 8 where ALSUV shows the best results.
0.C.4 Additional CFP-FP Results
Table 9: Additional SAR results of CFP-FP pairs compared with previous works with 800 pairs in total. We highlight our method with grey and best results are highlighted with bold.
Dataset CFP-FP
Test Encoder FaceNet MobFace ResNet50 ResNet100 Swin-S VGGNet Unseen AVG
Target Encoder Real Face 99.87 99.83 99.83 99.93 99.93 99.8 -
FaceNet NBNet 82 20.88 74.88 40.75 16.38 70.25 44.63
LatentMap 51.88 19.50 36.75 10.75 26.25 43.88 27.43
Genetic 60 23.25 41.62 10.62 29.25 42.12 29.37
GaussBlob 4.88 21 8 1.38 3.12 2.75 7.25
EigenFace 85 28.38 58.13 18.75 27 58.75 38.20
FaceTI 9.75 5.62 8.25 0.62 7.5 8.62 6.12
QEZOGE 93.75 39.12 81.62 62 65 83.62 66.27
Ours 94.5 47 87.38 74.12 77.12 90.12 75.15
MobileFaceNet NBNet 11.75 52 17.88 8 12.38 11.62 12.33
LatentMap 13.12 16.75 13 2.62 13.25 15 11.40
Genetic 18 35.38 16.5 5.38 18.12 16.88 14.98
GaussBlob 8.12 62.38 12.25 5.25 11.62 10.75 9.60
EigenFace 52.25 78.62 55.5 45.38 67.12 55.38 55.13
FaceTI 23.88 26.12 24.12 8.88 20.75 22.5 20.03
QEZOGE 45.25 74.62 50 39 57.75 49.12 48.22
Ours 85.25 82.88 87 86.75 87 86.88 86.58
ResNet50 NBNet 68.25 25.5 80.12 49.75 10.25 68.88 44.53
LatentMap 30.75 18.5 32.88 10.25 22.25 28.88 22.13
Genetic 38.62 24.25 54.37 11.62 28.25 39.75 28.5
GaussBlob 1.75 18.5 6.5 1 2.88 3.5 5.53
EigenFace 68.25 38.88 83.88 46.88 49 73.12 55.23
FaceTI 20.88 13.25 22.12 4.62 14.25 18.12 14.22
QEZOGE 79.12 42.25 89.12 61.88 67.5 80.38 66.23
Ours 92.62 66 94.38 85.75 86.62 90.88 84.37
ResNet100 NBNet 33.75 27.88 46.38 56 27.5 40.5 35.20
LatentMap 21.38 14.88 18.25 10.5 22.88 20.38 19.55
Genetic 27 21 22.88 24.5 28.12 25 24.8
GaussBlob 33.88 35.25 45.25 60.75 47.38 37.75 39.90
EigenFace 72 56 72.88 88.25 81 71.62 70.7
FaceTI 31.87 42.88 41.62 52.12 57.88 39.62 42.77
QEZOGE 56.62 41.5 60.38 88 69.5 61.5 57.9
Ours 92.5 74.25 92.75 93.62 90.12 90.38 88
Swin-S NBNet 26 29 37.12 31.75 32.75 29 30.57
LatentMap 16.12 14.37 14.37 5 21.62 13.75 12.72
Genetic 29 22.25 28.38 9.88 45.38 27 23.30
GaussBlob 4.62 29.25 7.12 4.62 17.25 4.5 10.02
EigenFace 26.88 39.38 36.5 25.5 55.38 30 31.65
FaceTI 46.5 47.88 53.25 43.5 63.75 50.5 48.33
QEZOGE 70.62 52 72.38 72.12 86.25 71.25 67.67
Ours 90.38 76.38 89.5 91.62 90.5 89.62 87.5
VGGNet19 NBNet 62.75 22.25 72.12 42.12 22.25 72.12 44.30
LatentMap 26.75 16.25 22.75 5.75 18 29 17.90
Genetic 46.88 23.75 44.25 11.88 27.5 57.38 30.85
GaussBlob 5.75 20.38 6.38 0.88 1.75 7.88 7.03
EigenFace 87 36.88 83.75 59.75 60 91.62 65.48
FaceTI 9.12 7.38 10.38 1.5 7.12 12.5 7.10
QEZOGE 82.12 43.5 81.25 62.38 68.12 90.75 67.47
Ours 93.5 61.12 91.38 84.88 85.12 93.5 83.20
We conduct additional SAR evaluation on CFP-FP dataset with 800 pairs of frontal-profile pairs which is the maximum amount of positive pairs from the 200 generated frontal images in Tab. 9. ALSUV shows the best performance comparing with previous baselines.
0.C.5 Additional Baseline Comparison
We additionally compare SAR results from LFW with GAFAR [31] which is a blackbox method, and a whitebox method [30] denoted as ICIP22. Our method surpasses both methods shown in Tab. 10.
Table 10: Comparing with additional baselines with LFW dataset.
FaceNet MobFaceNet ResNet50 Swin-S VGGNet AVG
GAFAR[31] 20.9 26.42 23.64 14.17 27.62 22.55
ICIP22[30] 59.22 82 64.98 75.6 70.17 70.68
Ours 99.36 96.16 99.56 99.93 99.56 98.91
0.C.6 ALSUV in Blackbox Scenario
ALSUV can be applied to blackbox scenario where gradients are not provided by combining with Zeroth-Order Gradeint Estimation(ZOGE). We compare with other blackbox methods in order to show the efficacy of ALSUV in blackbox scenario with results in Tab. 11 with averaged results in unseen scenario. For training setup, we increase the total iteration to 300 with weight decay of 0.1 at iteration 150 with Adam optimizer. For hyperparameters, we use $n=20$, $t_{avg}=3$, and $k_{top}=5$.
Table 11: Comparing SAR of blackbox baselines and ALSUV generated under blackbox scenario where ALSUV is trained with Zeroth-order gradient estimation testing on LFW dataset. Results are the average of unseen scenario.
Dataset NBNet Latent Map Genetic Gaussian Blob Eigenface FaceTI QEZOGE black-box ALSUV
LFW 22.21 11.40 27.21 9.97 57.21 18.73 69.49 82.38
@@ 0.D.1 Face Recognition Encoders
0.D.2 Dataset
The description of the dataset used for our experiments are summarized in this section. For LFW and AgeDB-30, we compare reconstructed samples with every other positive samples except the one used for reconstruction, resulting in 3,166 and 3,307 pairs respectively. CFP-FP consists of frontal and profile images; we reconstruct only the frontal images and follow the challenging frontal-profile verification protocol, resulting in 130 pairs in total for results in Tab. 1. We conduct additional experiments in Supp. 0.C.1 Tab. 9 where we use all possible frontal-profile possible pairs resulting in 800 pairs for additional results. For identification, we set up generated samples as probes and every image as the gallery, resulting in 13,233, 2,000, and 5,298 samples in the gallery consisting of 5,749, 500, and 388 identities for LFW, CFP-FP, and AgeDB-30 respectively. For CFP-FP, we use generated frontal images for probes and profile images for the gallery considering the original design of the dataset.